

********************************************************************
UARS WEEKLY STATUS REPORT - 23 JUNE 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. Level 2 processing is halted due to bugs in the OA_SSPP_ATT
service. Patches to the OA_SSPP_ATT and OA_SAT_ATT by CSC to the PAC and
we await their release. Once this happens, we will incorporate them in
our software and re-start level 2 processing.
***********
SUSIM
***********
Nothing received.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
At the present time, PEM energy deposition is not being generated.
Recent data configuration changes have affected PEM software causing multiple
failures. These failures are being reviewed by the PEM team. In order to
correct for new telemetry requirements, PEM will need to update processing
code with a patch delivery. The delivery patch will not include changes which
would allow the PEM software to execute on a DEC Alpha machine.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
HALOE continues successful operations with all engineering parameters
within nominal operating region. Science data results continue to be
excellent.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG attended the Pacor-A (Packet Processor Automation) UARS
Mission Elements Subsystem Desing Review on 22 June 2000. This
is a system being designed under CSOC development activities to
modify the Integrated Operations Architecture to enable cost savings.
This Pacor-A system will provide the Level-Zero processing and near-
real-time data, production data sets, quicklook data sets, and raw
data archiving for UARS data. The schedule indicates that start of
Pacor-A operations for UARS will begin in March 2001. Parallel
operation with the current system is planned for March, April, and
May, with the Operational Readiness Review planned for 7 May 2001.
The MPG also participated in a meeting with the UARS Mission
Director, the Project Scientist, FOT, and the GSFC Battery Branch
to discuss the condition of UARS batteries 2 and 3. It was
confirmed that these two batteries are continuing to function properly,
especially when we consider that they have been in orbit for over
eight years. There has been no change in battery conditions in the
last few weeks. (At least one cell has failed on Battery 1 and that
battery has been off-line since 1997.)
Processing of instrument Daily Activity Plans continues with no
problems. We are continuing to assist the Project Scientist in
scheduling of instrument operations. The schedule for instrument
operations during the next yaw period (4 July to 15 August) has
been distributed.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 10 June 2000 (Orbit 47807, GMT Day 162) through 16 June 2000
(Orbit 47911, GMT Day 168).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HRDI 164/0306 47839 Standby
HRDI 165/1730 47863 On
PEM 166/2058 47880 Day Only
The beta angle ranged between the angles of 13.7 degrees to 31.0 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
34.0 degrees on 20 June 2000 (DOY 172).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.6 degrees C. Additional battery performance data for each
battery on 09 June 2000 and 16 June 2000 are provided in Appendix D.
The clock error ranged from -11.4 to 11.2 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 04
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 23
% HRDI = 02
% PEM = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 01
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/145 - 2287.497008
- Transponder B: 00/152 - 2287.498920
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/160 show the observatory data loss to be 296 hours, 20
minutes, 7.713 seconds (an increase of 0 minutes 0.0 seconds since the last
report period). This is a 0.40 percent data loss which equals a 99.60 percent
data capture for the mission.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The testing of the new TMONs has been delayed due to a disk crash on the UTTS.
This disk was restored from tape but the backup tape used was from May due to
backup tape generation problems that were not fixed. Some coding changes and
simulation scenarios were lost in the process. This may require the FOT to
have to repeat some of its testing. Once the TMONs have been thoroughly
tested, they will be presented for review and approval before uplink.
The FOT has begun writing the online procedure for safehold should it happen
before the new TMONs are in place or during and OBC halt. This procedure will
be tested on the simulator and put online for immediate execution should an
anomalous condition take place.
MEETINGS
The FOT attended the UARS SERB meeting on 14 June 2000.
On 15 June 2000 the FOT attended the TDRS H Level 6 and Post Acceptance
Meeting. Testing of the new TDRS is tentatively scheduled to begin at the end
of August.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for July 04, 2000 (DOY 186). This maneuver
will be a Reverse to Forward Flight maneuver (Southern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 16 June 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
47810 162/0445 00-265 NBTR NBTR PLAYBACK
47816 162/1457 00-265 NBTR NOOPS D-162
47818 162/1825 00-267 NBTR NOOPS D-163
47826 163/0705 00-267 NBTR NOOPS D-163
47833 163/1827 00-268 NBTR NOOPS D-164
47839 164/0306 00-269 HRDI HRDI SAFEHOLD/PB OFF
47840 164/0450 00-268 NBTR NOOPS D-164
47846 164/1504 00-268 NBTR NOOPS D-164
47848 164/1734 00-270 NBTR NOOPS D-165
47855 165/0529 00-270 NBTR NOOPS D-165
47861 165/1507 00-270 NBTR NOOPS D-165
47863 165/1730 00-272 HRDI POWER ON; FULLTIME
47863 165/1832 00-271 NBTR NOOPS D-166
47870 166/0454 00-271 NBTR NOOPS D-166
47878 166/1733 00-271 NBTR NOOPS D-166
47878 166/1834 00-273 NBTR NOOPS D-167
47880 166/2058 00-274 PEM POWER ON/DAY ONLY
47885 167/0456 00-273 NBTR NOOPS D-167
47889 167/1138 00-275 OBC Clk Adj 0.00 to 20.6 msec/day
47891 167/1513 00-273 NBTR NOOPS D-167
47893 167/1837 00-276 NBTR NOOPS D-168
47899 168/0317 00-276 NBTR PLAYBACK
47900 168/0458 00-276 NBTR NOOPS D-168
47906 168/1515 00-277 OBC Clk Adj fr 20.6 to 0.00 msec/day
47906 168/1515 00-276 NBTR NOOPS D-168
47908 168/1739 00-278 NBTR NOOPS D-169
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 16 June 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRIFESURVTMP Y-HI IFE SURVIVAL TMP 47841-2
HRRADTEMP Y-HI IFE RADIATOR TMP 47841
HRPSEQBIMON Y-HI PSE QUIET BUS CUR MON 47899-02
HRPTSPBIMON Y-HI PTS PULSE BUS CUR MON 47899-01
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
-None this report period
______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 16 June 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
47889 167/11:57 (15 June 00) from 0.00 to 20.6 msec/day
47906 168/15:26 (16 June 00) from 20.6 to 0.00 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 16 June 2000
09 June 2000 - GMT Day 161 Beta = 13.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ----------- ------ ------- ------
1 +33.6/+11.2 -3.05/-4.62 +0.0/+0.0 15.7 0.000 0.0
2 +44.8/-22.4 +5.79/+4.51 +29.6/-16.0 24.3 1.030 20.6
3 +22.4/-67.2 +4.83/+3.23 +28.8/-15.6 24.2 1.027 20.2
16 June 2000 - GMT Day 168 Beta = 31.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ----------- ------- ------- ------
1 +39.2/+11.2 -2.42/-3.68 +0.0/+0.0 15.7 0.000 0.0
2 +28.0/-28.0 +7.09/+5.15 +28.8/-15.2 23.8 1.030 23.0
3 +22.4/-78.4 +5.47/+4.19 +28.4/-15.2 23.7 1.028 22.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 16 June 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-026 FOT 47829 163/1211 14170 Generic Late Acq.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-024 FOT 47742 157/1635 N/A TMON 1 trip
00-025 FOT 47766 159/0657 N/A TMONS 1 & 2 tripped
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiv 512
00-026 FOT see list June 00 14170 Generic Late Acq.
SEVNTFW00/24